OpenStudy (anonymous):

Let R be the planar region between the curves y^2 + (x-1)^2 = 4 and y^2 = 3(x-1) which contains the point (0, 0). 1.) Calculate the total area of ​​R. 2.) Determine the total length of the boundary of R.

OpenStudy (amistre64):

we got an interval to work with yet? or is this going to be one of those multivariate things

OpenStudy (amistre64):

y^2 + (x-1)^2 = 4 seems to be a circle centered at 1,0 of radius 2 y^2 = 3(x-1) 1/3 y^2 +1 = x parabola on its side with vertex at 1,0 i think

OpenStudy (anonymous):

i think y varies from (-sqrt(4-(x-1)^2) to (+sqrt(4+(x-1)^2)) and x to -1 to 1

OpenStudy (amistre64):



OpenStudy (amistre64):

think my parabola is a little off

OpenStudy (anonymous):

amistre is my upper and lower limits right?

OpenStudy (amistre64):

depends on which way youre integrating

OpenStudy (anonymous):

that + sqrt(4-(x-1)^2)

OpenStudy (amistre64):

 i was thinking of doing shells rotated about the x axis